OpenStudy (anonymous):

Evaluate. 2 • {390 – [5 • (10 – 4 ÷ 2)] + 15} • 2

OpenStudy (anonymous):

@Fairy733189

OpenStudy (fairy733189):

Do you want to try this one to see how you do?

OpenStudy (misssmartiez):

AS we said.

OpenStudy (fairy733189):

If you need help just ask

OpenStudy (anonymous):

yes

OpenStudy (fairy733189):

Great

OpenStudy (anonymous):

i'm lost

OpenStudy (fairy733189):

Show me what you have done so far

OpenStudy (misssmartiez):

1: 4 / 2 = 2. 2: 10 - 2 = 8. 3: 5 x 8 = 40 3: Anyway, it could be addition first, then subtraction, vice versa. I choose to do addition. 40 + 15 = 55. 390 - 55 = 335 4: easier way, 2 + 2 = 4. Now multiply that by 335. 1340.Still the same answer when you do it the hard way: 335 x 2 = 670 670 x 2 = 1340

OpenStudy (anonymous):

i got 1340n as the answer

OpenStudy (misssmartiez):

Correct.

OpenStudy (fairy733189):

Are you sure you understand though and you didn't just look at MissSmartiez?

OpenStudy (anonymous):

yes i just did it

OpenStudy (misssmartiez):

You have followed PEMDAS correctly. Normally some times questions can appear as: {30 x 4(5 + 2) - 4} - 10. Don't let the brackets fool you.

OpenStudy (fairy733189):

But do you understand it?

OpenStudy (anonymous):

i know

OpenStudy (fairy733189):

Great!!!

OpenStudy (fairy733189):

Do you have any more problems you need help with?

OpenStudy (misssmartiez):

You do PARENTHESIS first before ANY following bracket, then, you do the bracket, then outside the bracket. Pemdas, P - parenthesis E - exponent (skip if not present) M - multiplication D - division A - addition S - subtraction. Anymore questions?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i failed it

OpenStudy (fairy733189):

What do you mean?

OpenStudy (misssmartiez):

How so ^ What did you possibly get wrong ;)?

OpenStudy (anonymous):

idk

OpenStudy (fairy733189):

k bye

OpenStudy (misssmartiez):

Hmm, if you still are confused, tag me and we will go over it and some examples.